# gRNAde: Geometric Deep Learning for 3D RNA inverse design

### v4

@@ 4.3 Zero-shot ranking of RNA fitness landscape
Lastly, we explored the use of gRNAde as an zero-shot ranker of mutants in RNA engineering campaigns. Given the backbone structure of a wild type RNA of interest as well as a candidate set of mutant sequences, we can compute gRNAde’s perplexity of whether a given sequence folds into the backbone structure. Perplexity is inversely related to the likelihood of a sequence conditioned on a structure, as described in Section 2.4. We can then rank sequences based on how ‘compatible’ they are with the backbone structure in order to select a subset to be experimentally validated in wet labs.
Retrospective analysis on ribozyme fitness landscape. A recent study by McRae et al. (2024) determined a cryo-EM structure of an RNA polymerase ribozyme at 5Å resolution [^6] This RNA was not present in gRNAde’s training data, which contains structures at $\leq$4.0Å resolution., along with a fitness landscape of $\sim$75,000 mutant sequences for the catalytic subunit 5TU. We design a retrospective study using this data of (sequence, fitness value) pairs where we simulate an RNA engineering campaign with the aim of improving fitness over the wild type sequence.
@@
gRNAde outperforms random baselines in low design budget scenarios. Figure 7 illustrates the results of our retrospective study. At low design budgets of up to hundreds of sequences, which are relevant in the case of a low throughput fitness screening assay, gRNAde outperforms all random baselines in terms of the maximum change in fitness over the wild type. The top 10 mutants as ranked by gRNAde contains a sequence with 4-fold improved fitness, while the top 200 leads to a 5-fold improvement777 As a caveat, the fitness assays from McRae et al. (2024) used for creating the landscape have inherent noise and cannot easily differentiate between mutants of similar activity. . Beyond design budgets of thousands, random selection from double mutants starts outperforming gRNAde.
Perspective. Overall, it is promising that gRNAde’s perplexity is somewhat correlated with experimental fitness measurements out-of-the-box (zero-shot) and can be a useful ranker of mutant fitness in our retrospective study. In realistic design scenarios, improvements could likely be obtained by fine-tuning gRNAde on a low amount of experimental fitness data. For example, latent features from gRNAde may be finetuned or used as input to a prediction head with supervised learning on fitness landscape data. This study acts as a sanity check before committing to wet lab validation of gRNAde designs. We see random mutagenesis and directed evolution-based approaches as complementary to de-novo design and inverse folding approaches like gRNAde. Random mutagenesis can be thought of as local exploration around a wild type sequence, optimising fitness within an ‘island’ of activity. Structure-based design approaches are akin to global jumps in sequence space, with the potential to find new islands further away from the wild type (Huang et al., 2016).
Figure 7: Retrospective study of gRNAde for ranking ribozyme mutant fitness. Using the backbone structure and mutational fitness landscape data from an RNA polymerase ribozyme (McRae et al., 2024), we retrospectively analyse how well we can rank variants at multiple design budgets using random selection vs. gRNAde’s perplexity for mutant sequences conditioned on the backbone structure. Note that gRNAde is used zero-shot here, i.e. it was not fine-tuned on any assay data. For stochastic strategies, bars indicate median values, and error bars indicate the interquartile range estimated from 10,000 simulations per strategy and design budget. At low throughput design budgets of up to $\sim$500 sequences, selecting mutants using gRNAde outperforms random baselines in terms of the expected maximum improvement in fitness over the wild type. In particular, gRNAde performs better than single site saturation mutagenesis, even when all single mutants are explored (449 mutants for the RNA polymerase ribozyme in McRae et al. (2024)).
@@ 5 Conclusion
We introduce gRNAde, a geometric deep learning pipeline for RNA sequence design conditioned on one or more 3D backbone structures. gRNAde is superior to the physically based Rosetta for 3D RNA inverse folding in terms of performance, inference speed, and ease of use. Additionally, gRNAde enables explicit multi-state design for structurally flexible RNAs which was previously not possible with Rosetta. To the best of our knowledge, gRNAde is also the first geometric deep learning architecture for multi-state biomolecule representation learning; the model is generic and can be repurposed for other learning tasks on conformational ensembles, including multi-state protein design. Key limitations of gRNAde and avenues for future development include the lack of support for multiple interacting RNA chains, or accounting for biomolecular interactions of RNAs with proteins, small molecules, and other ligands. More advanced user configurations, such as negative design against undesired conformations or modulating conformational propensities are also not currently possible. Finally, we are hopeful that advances in RNA structure determination and computationally assisted cryo-EM (Kappel et al., 2020; Bonilla and Kieft, 2022) will further increase the amount of RNA structures available for training geometric deep learning models in the future.
@@ Acknowledgements
Appendix A Related Work
@@
Appendix B FAQs on using gRNAde
@@
How to prioritise or chose amongst designed sequences? We have currently provided 3 evaluation metrics: native sequence recovery, secondary structure self-consistency score and perplexity, towards this end. We suspect that recovery may not be the ideal choice, except for design scenarios where we require certain regions of the RNA to be conserved or native-like. Self-consistency score may be a more holistic evaluation metric as it accounts for alternative base pairings leading to similar structures, but may inherit the limitations of the secondary structure prediction method used as part of its computation. In realistic design scenarios, we can pair gRNAde with another machine learning model (an ‘oracle’) for ranking or predicting the suitability of designed sequences for the objective (for instance, binding affinity or some other notion of fitness). We hope to conduct further experimental validation of gRNAde designs in the wet lab in order to better understand these tradeoffs.
Appendix C Additional Figures and Tables
Figure 8: In-silico evaluation metrics for gRNAde.

### v5

@@ 4.3 Zero-shot ranking of RNA fitness landscape
Lastly, we explored the use of gRNAde as a zero-shot ranker of mutants in RNA engineering campaigns. Given the backbone structure of a wild type RNA of interest as well as a candidate set of mutant sequences, we can compute gRNAde’s perplexity of whether a given sequence folds into the backbone structure. Perplexity is inversely related to the likelihood of a sequence conditioned on a structure, as described in Section 2.4. We can then rank sequences based on how ‘compatible’ they are with the backbone structure in order to select a subset to be experimentally validated in wet labs.
Retrospective analysis on ribozyme fitness landscape. A recent study by McRae et al. (2024) determined a cryo-EM structure of a dimeric RNA polymerase ribozyme at 5Å resolution [^4] This RNA was not present in gRNAde’s training data, which contains structures at $\leq$4.0Å resolution., along with fitness landscapes of $\sim$75K mutants for the catalytic subunit 5TU and $\sim$48K mutants for the scaffolding subunit t1. We design a retrospective study using this data of (sequence, fitness value) pairs where we simulate an RNA engineering campaign with the aim of improving catalytic subunit fitness over the wild type 5TU sequence.
@@
gRNAde outperforms random baselines in low design budget scenarios. Figure 7 illustrates the results of our retrospective study. At low design budgets of up to hundreds of sequences, which are relevant in the case of a low throughput fitness screening assay, gRNAde outperforms all random baselines in terms of the maximum change in fitness over the wild type. The top 10 mutants as ranked by gRNAde contain a sequence with 4-fold improved fitness, while the top 200 leads to a 5-fold improvement555 As a caveat, the fitness assays from McRae et al. (2024) used for creating the landscape have inherent noise and cannot easily differentiate between mutants of similar activity.. Note that gRNAde is used zero-shot here, i.e. it was not fine-tuned on any assay data.
Perspective. Overall, it is promising that gRNAde’s perplexity correlates with experimental fitness measurements out-of-the-box (zero-shot) and can be a useful ranker of mutant fitness in our retrospective study. In realistic design scenarios, improvements could likely be obtained by fine-tuning gRNAde on a low amount of experimental fitness data. For example, latent features from gRNAde may be finetuned or used as input to a prediction head with supervised learning on fitness landscape data. This study acts as a sanity check before committing to wet lab validation of gRNAde designs. We see random mutagenesis and directed evolution-based approaches as complementary to de-novo design and inverse folding approaches like gRNAde. Random mutagenesis can be thought of as local exploration around a wild type sequence, optimising fitness within an ‘island’ of activity. Structure-based design approaches are akin to global jumps in sequence space, with the potential to find new islands further away from the wild type (Huang et al., 2016).
Figure 7: Retrospective study of gRNAde for ranking ribozyme mutant fitness. Using the backbone structure and mutational fitness landscape data from an RNA polymerase ribozyme (McRae et al., 2024), we retrospectively analyse how well we can rank variants at multiple design budgets using random selection vs. gRNAde’s perplexity for mutant sequences conditioned on the backbone structure (catalytic subunit 5TU). Note that gRNAde is used zero-shot here, i.e. it was not fine-tuned on any assay data. For stochastic strategies, bars indicate median values, and error bars indicate the interquartile range estimated from 10,000 simulations per strategy and design budget. At low throughput design budgets of up to $\sim$500 sequences, selecting mutants using gRNAde outperforms random baselines in terms of the expected maximum improvement in fitness over the wild type. In particular, gRNAde performs better than single site saturation mutagenesis, even when all single mutants are explored (total of 449 single mutants, 10,493 double mutants for the catalytic subunit 5TU in McRae et al. (2024)). See Appendix Figure 12 for results on scaffolding subunit t1.
@@ 5 Conclusion
We introduce gRNAde, a geometric deep learning pipeline for RNA sequence design conditioned on one or more 3D backbone structures. gRNAde is superior to the physically based Rosetta for 3D RNA inverse folding in terms of performance, inference speed, and ease of use. Further, gRNAde enables explicit multi-state design for structurally flexible RNAs which was previously not possible with Rosetta. gRNAde’s perplexity correlates with native sequence and structural recovery, and can be used for zero-shot ranking of mutants in RNA engineering campaigns. To the best of our knowledge, gRNAde is also the first geometric deep learning architecture for multi-state biomolecule representation learning; the model is generic and can be repurposed for other learning tasks on conformational ensembles, including multi-state protein design.
Key avenues for future development of gRNAde include supporting multiple interacting chains, accounting for partner molecules with RNAs, and supporting negative design against undesired conformations. We discuss practical tradeoffs to using gRNAde in real-world RNA design scenarios in Appendix C, including limitations due to the current state of 3D RNA structure prediction tools. Finally, we are hopeful that advances in RNA structure determination and computationally assisted cryo-EM (Kappel et al., 2020; Bonilla and Kieft, 2022) will further increase the amount of RNA structures available for training geometric deep learning models in the future.
@@ Acknowledgements
Appendix A 3D Visualisation of gRNAde Designs
(a) self-consistency TM-score
(b) self-consistency GDT_TS
Figure 8: 3D self-consistency scores for 3 representative RNAs from Das et al. (2010). We use RhoFold to ‘forward fold’ 100 designs sampled at temperature = 0.5 and plot self-consistency TM-score and GDT_TS. Each dot corresponds to one designed sequence and is coloured by gRNAde’s perplexity (normalised per RNA). Designs with lower relative perplexity generally have higher 3D self-consistency and can be considered more ‘designable’. Dotted lines represent TM-score and GDT_TS thresholds of 0.45 and 0.50, repsectively. Pairs of structures scoring higher than the threshold correspond to roughly the same fold.
Design 1:
GGCAAGUAAUCCCUACGCUAUG
GGUAGGGAGUCUCAGCAGUGAC
CCGUAAAGUUACUACCUUGCCC
perplexity: 1.3097
recovery: 0.5909 (27 edits)
sc2D = 0.9227
scRMSD = 1.3839
scTM = 0.8309
scGDT = 0.8295
Design 2:
CGGUGGUAAGCCCAACGCUAGG
GGUUGGGCGUCUCAGCACAGUC
CCGUAAAGAUUGUACCCACCGG
perplexity: 1.3815
recovery: 0.4091 (37 edits)
sc2D = 0.9227
scRMSD = 2.1249
scTM = 0.6874
scGDT = 0.6780
Design 3:
AGCAAGUAAUGCCAUCGCUAUG
GGAUGGUAGUGUCAGCACUGAC
CCUUAAAGUUAGUACCUUGCUU
perplexity: 1.4247
recovery: 0.5152 (30 edits)
sc2D = 0.9227
scRMSD = 3.2131
scTM = 0.5118
scGDT = 0.5265
Figure 9: Cherry-picked designs for Guanine riboswitch aptamer (PDB: 4FE5, sequence: GGACAUAUAAUCGCGUGGAUAUGGCACGCAAGUUUCUACCGGGCACCGUAAAUGUCCGACUAUGUCC).
Design 1:
GGGGCUCCGGCGACGCAGUCGAAAG
CCCAGCAGUACCAAGCCUCAGGGGA
AACUUUGAGGUGGCCUAACAAAGGA
UACGGUAAUAAGCUGCGGGAAAAGG
UUGUAAGCCGGAGCGAAGACCUAAG
GCACCGCUUUUGGCGGUGCUAUGGU
UGAAGUUAA
perplexity: 1.2462
recovery: 0.7170 (44 edits)
sc2D = 0.8301
scRMSD = 5.4562
scTM = 0.6481
scGDT = 0.4465
Design 2:
GGGGUACCGGCGACGCAGUCGAAUG
CCCUGUGGUACCAAGCCCCGGGGGA
AACUUCGGGGUGGCCUUACCAAGGA
CACGGUAAUAAGCCACGGGAAAUGG
UUGUAAGCCGGUCCGAAGCCCUAAG
GCCGCGCUUUGGGCGCGGCUAUGGG
UGAAGGCAA
perplexity: 1.3273
recovery: 0.6226 (58 edits)
sc2D = 0.6896
scRMSD = 6.7239
scTM = 0.6300
scGDT = 0.4513
Design 3:
GAGGCCACGGCAACGCAGUCUAACG
CCCUGUGGUACCAAGUCUUAGGAGA
AAUUUUAAGAUGGCCUAAUAAAGGA
UAUGGUAAUAAGCCACGGGAAAAGG
UUGUAAGACGUGACGAAGUCCUAAG
GCCACAGUUUUGCUGUGGCUAUGGA
UGGAGUACA
perplexity: 1.3204
recovery: 0.7044 (45 edits)
sc2D = 0.7922
scRMSD = 8.8211
scTM = 0.4582
scGDT = 0.2909
Figure 10: Cherry-picked designs for Tetrahymena Ribozyme P4-P6 domain (PDB: 2R8S, sequence: GGAAUUGCGGGAAAGGGGUCAACAGCCGUUCAGUACCAAGUCUCAGGGGAAACUUUGAGAUGGCCUUGCAAAGGGU
AUGGUAAUAAGCUGACGGACAUGGUCCUAACACGCAGCCAAGUCCUAAGUCAACAGAUCUUCUGUUGAUAUGGAUGCAGUUCA
Design 1:
GUCAAACGCAGCCGAAA
GCGCGAUAGUCCCAGGAA
perplexity = 1.6237
recovery = 0.4571 (16 edits)
sc2D = -0.0074
scRMSD = 3.9505
scTM = 0.2597
scGDT = 0.4786
Design 2:
GGCAAACGCGGCCGAAA
GCGCGUGAGUCCCCGGAC
perplexity = 1.6630
recovery = 0.4857 (16 edits)
sc2D = -0.0099
scRMSD = 3.3549
scTM = 0.2526
scGDT = 0.5000
Design 3:
CGUAGUCGGAGCCGAAG
GGCCGUUAGUCCCAGGAG
perplexity = 1.7020
recovery = 0.4000 (17 edits)
sc2D = 0.4035
scRMSD = 16.4102
scTM = 0.0319
scGDT = 0.0571
Figure 11: Cherry-picked designs for Vitamin B12 binding aptamer (PDB: 1ET4, sequence: GGAACCGGUGCGCAUAACCACCUCAGUGCGAGCAA).
Appendix B Related Work
@@
Appendix C FAQs on using gRNAde
@@
How to prioritise or chose amongst designed sequences? We have currently provided 3 types of evaluation metrics: native sequence recovery, structural self-consistency scores and perplexity, towards this end. We suspect that recovery may not be the ideal choice, except for design scenarios where we require certain regions of the RNA sequence to be conserved or native-like. Self-consistency scores may provide an overall more holistic evaluation metric as they accounts for alternative base pairings which still lead to similar structures as well as better capture the recovery of structural motifs responsible for functionality. However, structural self-consistency scores inherit the limitations of the structure prediction methods used as part of their computation. For instance, computing the self-consistency score between an RNA backbone and its own native sequence provides an upper bounds on the maximum score that designs can obtain under a given structure prediction method. Lastly, gRNAde’s perplexity estimates the likelihood of a sequence given a backbone and can be useful for ranking designs and mutants in RNA engineering campaigns (especially for design scenarios where structure prediction tools are not performant).
In real-world design scenarios, we can pair gRNAde with another machine learning model (an ‘oracle’) for ranking or predicting the suitability of designed sequences for the objective (for instance, binding affinity or some other notion of fitness). We hope to conduct further experimental validation of gRNAde designs in the wet lab in order to better understand these tradeoffs.
Why not average single-state logits over multiple states for multi-state design? ProteinMPNN (Dauparas et al., 2022) proposes to average logits from multiple backbones for multi-state protein design. Here is a simple example to highlight issues with such an approach: Consider two states A and B, and choice of labels X, Y, and Z. For state A: X, Y, Z are assigned probabilities 75%, 20%, 5%. For state B: X, Y, Z are assigned probabilities 5%, 20%, 75%. Logically, label Y is the only one that is compatible with both states. However, averaging the probabilities would lead to label X or Z being more likely to be sampled in designs. As an alternative, gRNAde is based on multi-state GNNs which can take as input one or more backbone structures and generate sequences conditioned on the conformational ensemble directly.
Appendix D Ablation Study
Table 1 presents an ablation study as well as aggregated benchmark for various configurations of gRNAde. Key takeaways are highlighted below. Note that all results in the main paper are reported for models trained on the maximum length of 5000 nucleotides using autoregressive decoding and rotation-equivariant GNN layers, as this lead to the lowest perplexity values.
[…]
Self-consistency metrics Max. Max. train Perplexity Native seq. 2D – EternaFold 3D – RhoFold Split #states Model GNN length ($\downarrow$) recovery ($\uparrow$) scMCC ($\uparrow$) scRMSD ($\downarrow$) scTM-score ($\uparrow$) scGDT_TS ($\uparrow$) Single-state split 1 AR Equiv 500 1.77$\pm$0.07 0.438$\pm$0.01 0.624$\pm$0.07 13.01$\pm$1.18 (0.5%) 0.21$\pm$0.0 (14.3%) 0.22$\pm$0.0 (12.7%) 1 AR Equiv 1000 1.73$\pm$0.08 0.453$\pm$0.01 0.648$\pm$0.01 13.10$\pm$0.58 (1.0%) 0.20$\pm$0.0 (10.8%) 0.21$\pm$0.0 (10.6%) 1 AR Equiv 2500 1.41$\pm$0.01 0.493$\pm$0.01 0.633$\pm$0.03 11.76$\pm$0.91 (1.4%) 0.27$\pm$0.0 (28.8%) 0.27$\pm$0.0 (28.0%) 1 AR Equiv 5000 1.29$\pm$0.02 0.530$\pm$0.01 0.585$\pm$0.03 11.70$\pm$0.56 (1.3%) 0.26$\pm$0.0 (24.8%) 0.25$\pm$0.0 (20.1%) 1 AR Inv 5000 1.32$\pm$0.04 0.549$\pm$0.00 0.612$\pm$0.02 11.50$\pm$0.64 (1.9%) 0.28$\pm$0.0 (32.1%) 0.28$\pm$0.0 (26.2%) 1 NAR Inv 5000 1.54$\pm$0.04 0.571$\pm$0.00 0.430$\pm$0.02 14.26$\pm$0.51 (1.3%) 0.19$\pm$0.0 (15.9%) 0.18$\pm$0.0 (12.7%) 1 NAR Equiv 5000 1.46$\pm$0.06 0.584$\pm$0.00 0.473$\pm$0.02 13.04$\pm$0.88 (1.3%) 0.23$\pm$0.0 (24.0%) 0.22$\pm$0.0 (17.9%) 3 AR Equiv 5000 1.23$\pm$0.05 0.539$\pm$0.01 0.620$\pm$0.01 11.47$\pm$1.05 (2.5%) 0.28$\pm$0.0 (31.4%) 0.28$\pm$0.0 (27.2%) 5 AR Equiv 5000 1.25$\pm$0.01 0.539$\pm$0.02 0.596$\pm$0.03 11.90$\pm$1.00 (2.9%) 0.27$\pm$0.0 (31.6%) 0.26$\pm$0.0 (26.4%) Groundtruth sequence prediction baseline: - 1.000$\pm$0.00 0.686$\pm$0.00 5.23$\pm$0.07 (27.9%) 0.56$\pm$0.0 (68.7%) 0.55$\pm$0.0 (68.7%) Random sequence prediction baseline: - 0.251$\pm$0.00 0.012$\pm$0.00 24.40$\pm$0.34 (0.0%) 0.04$\pm$0.0 (0.0%) 0.02$\pm$0.0 (0.0%) ViennaRNA 2D-only baseline: - 0.259$\pm$0.00 0.611$\pm$0.00 20.34$\pm$0.10 (0.0%) 0.07$\pm$0.0 (0.6%) 0.07$\pm$0.0 (1.1%) Multi-state split 1 AR Equiv 500 1.87$\pm$0.06 0.445$\pm$0.01 0.603$\pm$0.03 13.08$\pm$0.20 (3.5%) 0.10$\pm$0.0 (1.2%) 0.25$\pm$0.0 (20.7%) 1 AR Equiv 1000 1.84$\pm$0.01 0.447$\pm$0.01 0.580$\pm$0.01 13.02$\pm$0.56 (2.3%) 0.09$\pm$0.0 (0.9%) 0.25$\pm$0.0 (20.4%) 1 AR Equiv 2500 1.73$\pm$0.04 0.480$\pm$0.02 0.567$\pm$0.01 12.83$\pm$0.05 (3.4%) 0.10$\pm$0.0 (1.9%) 0.26$\pm$0.0 (21.2%) 1 AR Equiv 5000 1.68$\pm$0.03 0.455$\pm$0.01 0.569$\pm$0.02 12.88$\pm$0.20 (4.1%) 0.11$\pm$0.0 (1.6%) 0.26$\pm$0.0 (22.6%) 1 AR Inv 5000 1.72$\pm$0.01 0.463$\pm$0.01 0.559$\pm$0.03 13.09$\pm$0.27 (4.1%) 0.10$\pm$0.0 (2.2%) 0.27$\pm$0.0 (23.0%) 1 NAR Inv 5000 2.01$\pm$0.04 0.457$\pm$0.01 0.461$\pm$0.01 14.06$\pm$0.23 (3.2%) 0.08$\pm$0.0 (1.7%) 0.23$\pm$0.0 (16.5%) 1 NAR Equiv 5000 1.89$\pm$0.06 0.432$\pm$0.01 0.423$\pm$0.01 13.63$\pm$0.27 (3.6%) 0.09$\pm$0.0 (1.2%) 0.24$\pm$0.0 (18.3%) 3 AR Equiv 5000 1.60$\pm$0.03 0.467$\pm$0.03 0.561$\pm$0.03 13.31$\pm$0.38 (3.4%) 0.10$\pm$0.0 (2.6%) 0.24$\pm$0.0 (19.0%) 5 AR Equiv 5000 1.55$\pm$0.04 0.473$\pm$0.01 0.549$\pm$0.03 13.48$\pm$0.79 (3.3%) 0.10$\pm$0.0 (3.0%) 0.24$\pm$0.0 (20.2%) Groundtruth sequence prediction baseline: - 1.000$\pm$0.00 0.570$\pm$0.01 9.78$\pm$0.13 (10.3%) 0.16$\pm$0.0 (11.7%) 0.36$\pm$0.0 (36.7%) Random sequence prediction baseline: - 0.249$\pm$0.00 0.128$\pm$0.00 21.15$\pm$0.21 (0.9%) 0.02$\pm$0.0 (0.0%) 0.09$\pm$0.0 (3.3%) ViennaRNA 2D-only baseline: - 0.258$\pm$0.00 0.601$\pm$0.00 15.47$\pm$0.20 (2.4%) 0.05$\pm$0.0 (0.2%) 0.19$\pm$0.0 (15.2%) All data 1 AR Equiv 5000 1.23$\pm$0.01 0.733$\pm$0.00 0.627$\pm$0.02 8.10$\pm$0.28 (20.7%) 0.42$\pm$0.0 (46.1%) 0.41$\pm$0.0 (43.0%) 2 AR Equiv 5000 1.21$\pm$0.01 0.783$\pm$0.01 0.629$\pm$0.03 8.40$\pm$0.09 (19.1%) 0.42$\pm$0.0 (47.8%) 0.41$\pm$0.0 (41.7%) 3 AR Equiv 5000 1.19$\pm$0.01 0.787$\pm$0.01 0.606$\pm$0.02 7.88$\pm$0.68 (20.5%) 0.43$\pm$0.0 (47.4%) 0.42$\pm$0.0 (44.0%) 5 AR Equiv 5000 1.15$\pm$0.01 0.811$\pm$0.01 0.617$\pm$0.02 7.51$\pm$0.30 (20.7%) 0.45$\pm$0.0 (50.2%) 0.44$\pm$0.0 (46.7%)
Max. train RNA length Limiting the maximum length of RNAs used for training can be seen as ablating the use of ribosomal RNA families (which are thousands of nucleotides long and form complexes with specialised ribosomal proteins). We find that training on only short RNAs fewer than 1000s of nucleotides leads to worse sequence recovery and 3D self-consistency scores, even though it improves 2D self-consistency across both evaluation splits. This suggests that tertiary interactions learnt from ribosomal RNAs can generalise to other RNA families to some extent (large ribosomal RNAs were excluded from test sets).
GNN We ablated whether the internal representations of the GVP-GNN are rotation invariant or equivariant. Equivariant GNNs are theoretically more expressive (Joshi et al., 2023) and we do find them more capable at fitting the training distribution (as shown by lower perplexity). However, we do not find significant differences in terms of other performance metrics across different GNN layers.
Model ‘AR’ implies autoregressive decoding (described in Section 2.3, uses 4 encoder and 4 decoder layers), while ‘NAR’ implies non-autoregressive, one-shot decoding using an MLP (uses 8 encoder layers). Across both evaluation splits, AR models show significantly higher self-consistency scores than NAR, even though NAR lead to higher sequence recovery. AR is more expressive and can condition predictions at each decoding step on past predictions, while one-shot NAR samples from independent probability distributions for each nucleotide. Thus, AR is a better inductive bias for predicting base pairing and base stacking interactions that are drivers of RNA structure (Vicens and Kieft, 2022). For instance, G-C and A-U pairs can often be swapped for one another, but non-autoregressive decoding does not capture such paired constraints.
Max. #states We evaluate the impact of increasing the maximum number of states as input to gRNAde. Multi-state models marginally improve native sequence recovery as well as structural self-consistency scores over an equivalent single state variant, even for the single-state benchmark where the multi-state model is being used with only one state as input. This suggests that seeing multiple states during training can be useful for gRNAde’s performance even for single-state design tasks.
Non-learnt baselines. We report the performance of two non-learnt baselines to contextualise gRNAde’s performance: for each test sample, simply predicting the groundtruth sequence back and predicting a random sequence. Structural self-consistency scores for the Groundtruth baseline provides a rough upper bounds on the maximum score that any gRNAde designs can theoretically obtain given the current state of 2D/3D structure predictors being used. gRNAde always performs better than the random baseline and often reaches 2D self-consistency scores close to the upper bound. Both 2D and 3D self-consistency scores are inherently limited by the performance of the structure prediction methods used.
2D inverse folding baseline. We additionally report results for ViennaRNA’s 2D-only inverse folding method to further demonstrate the utility of 3D inverse folding. ViennaRNA has improved 2D self-consistency scores over gRNAde but fails to capture tertiary interactions in its designs, as evident by poor recovery and 3D self-consistency scores similar to the random baseline.
Split. Single- and multi-state splits are described in Section 3; the multi-state split is relatively harder than the single-state split based on overall reduced performance for all baselines and models. Models trained on ‘All data’ use all RNASolo samples for training, solely for the purpose of releasing the best possible gRNAde checkpoints for real-world usage. Evaluation metrics for ‘All data’ are reported on the single-state test set.
Appendix E Additional Results